OpenStudy (anonymous):

find domain and range of 1-√x

OpenStudy (anonymous):

Let: y = 1 - \(\sqrt{x}\) now tell me what are the values that x cannot take.. x cannot be negative..

OpenStudy (anonymous):

So x must be greater than or equal to zero.. because root of negative number is not defined.. So: \[x \ge 0\] This is the domain..

OpenStudy (anonymous):

Getting? @dorkkk

OpenStudy (anonymous):

how do i get the range? O:

OpenStudy (anonymous):

Can you find x in terms of y??

OpenStudy (anonymous):

what does range mean exactly?

OpenStudy (anonymous):

Subtract 1 both the sides: y - 1 = - \(\sqrt{x}\) Square both the sides: \[x=(y-1)^2\]

OpenStudy (anonymous):

Range you can think it as output.. See you have given x as input and you have got output as y or f(x).. So x is Domain and y or f(x) is range.. Getting??

OpenStudy (anonymous):

So in: \[x = (y-1)^2\] Here you can give any values to the y the equation will be determined in all the case.. So Range is All Reals..

OpenStudy (anonymous):

ahh, okay i get it. thanks!

OpenStudy (anonymous):

would i write all real numbers as (∞,∞)?